# Online Bayesian Experimental Design for Partially Observable Dynamical Systems

### v1

@@ 5.2 Moving source location
with cardioid directivity $D(\delta)=\big(\frac{1+d\cos\delta}{1+d}\big)^{k}$ ($d\in[0,1)$), background $b>0$, saturation $m>0$, directional strength exponent $k>0$ and sensor strengths $\alpha_{j}\geq 0$. The design $\boldsymbol{\xi}_{t}$ affects only the observation model.
Experimental setup.
We simulate a moving source over a horizon of T=50T\!=\!50 steps starting at position $\boldsymbol{x}_{0}\!=\!(0,0,0)$. The source parameters to be inferred are the horizontal and vertical velocities, $\boldsymbol{\theta}=(v_{x},v_{y})$, while the angular velocity $v_{\phi}$ and all other observation parameters ($b$, $m$, $d$, $k$, $\alpha_{j}$, $\sigma^{2}$) are treated as known. We deploy J=2J\!=\!2 fixed sensors located at $\boldsymbol{s}_{1}\!=\!(0,3)$ and $\boldsymbol{s}_{2}\!=\!(3,0)$, each with orientation design variables $\boldsymbol{\xi}_{t}\!=\!(\xi_{t,1},\xi_{t,2})\in[-\pi,\pi)^{2}$. Figure 3 illustrates the setup and one example trajectory. Inference is performed with a NPF using M=N=300M\!=\!N\!=\!300 particles, and design optimization at each step uses K=300K\!=\!300 stochastic gradient updates. Optimization is carried out with Adam (DBLP:journals/corr/KingmaB14). Additional implementation details and parameter values are given in Appendix LABEL:supp-ap:source.
Figure 3: Example trajectory of the moving source for t=50t\!=\!50 (blue line with markers) together with fixed sensor locations (red triangles).
Experimental results.
Table 1 reports the total EIG (TEIG) for our sequential method compared against the random and static baselines. Once again, our approach achieves the highest TEIG across all horizons, confirming the benefits of adapting designs online in partially observable settings. Figure 4 illustrates these results more clearly by showing boxplots of TEIG differences relative to each baseline. The improvements accumulate steadily over time as information gain compounds. As in the SIR experiment, results for the static method are truncated at larger horizons due to computational infeasibility of offline optimization in high-dimensional design spaces.
Figure 4: Moving source location: differences in TEIG (ours minus baseline). Boxplots (median and interquartiles) compare BAD-PODS with random designs (left, green) and static BED (right, orange). Higher values indicate better performance for BAD-PODS.
Finally, we assess the quality of the selected sensor orientations in terms of their pointing error, i.e., the angular deviation between a sensor’s chosen orientation and the true bearing to the source. Figure 5 shows boxplots of pointing error distributions at selected time steps across seeds and sensors. Random designs exhibit wide variability, often pointing far away from the target. The static baseline is more stable, but it still accumulates errors over time, as it cannot adapt to the evolving state of the system. In contrast, our sequential method consistently achieves much smaller pointing errors, demonstrating its ability to align sensors toward the target more frequently and with greater accuracy.
Figure 5: Pointing error (in degrees) of sensor orientations relative to the target at selected time steps. Boxplots (median and interquartiles) show the distribution across seeds and sensors for our method, random design, and static BED approach. Lower values correspond to higher orientation accuracy.
6 CONCLUSIONS
We introduced an online Bayesian adaptive design (BAD) framework for partially observable dynamical systems, modeled as state-space models (SSMs). Our approach combines two key ingredients: (i) new estimators of the expected information gain and its gradient that explicitly account for latent states, and (ii) an nested particle filter (NPF)-based construction that reuses state–parameter particles to evaluate these quantities online. This yields an algorithm whose cost grows only linearly with time, avoids reprocessing past data, and remains asymptotically consistent under mild assumptions. The framework supports continuous design spaces and stochastic-gradient optimization, making it broadly applicable. Through experiments on a multi-group epidemiological model and a dynamical source–tracking problem, we demonstrated that online adaptation significantly improves information efficiency over static Bayesian experimental design (BED) or random design baselines.

### v2

@@ F.1 Model Description
for $j=1,\ldots,J$, where $b>0$ is a background level, $m>0$ a saturation constant, $\alpha_{j}\geq 0$ a per-sensor strength, and
$D(\delta)\;=\;\bigg(\frac{1+d\cos\delta}{1+d}\bigg)^{k},\qquad d\in[0,1),\quad k>1,$
is a cardioid directivity function that favors alignment between the sensor orientation and the bearing to the source, while ensuring that $\mu_{t,j}>0$ for all configurations. The design vector $\boldsymbol{\xi}_{t}$ enters the observation model (20) only through the angular offset $\Delta_{t,j}$.
F.2 Simulation Setup
We use $J=2$ fixed sensors at $\boldsymbol{s}_{1}\!=\!(3,0)^{\top}$ and $\boldsymbol{s}_{2}\!=\!(0,3)^{\top}$. Per–sensor strengths are $\alpha_{j}\!=\!5$ (for j=1,2j\!=\!1,2), with background level b=0.1b\!=\!0.1 and saturation constant m=0.1m\!=\!0.1. The angular directivity $D(\delta)$ is set with d=1d\!=\!1 and k=4k\!=\!4. Observations are log–intensities with i.i.d. Gaussian noise $\sigma^{2}\!=\!0.1$, hence $R\!=\!\sigma^{2}\mathbf{I}_{J}$. Designs are initialized uniformly as $\boldsymbol{\xi}_{t}\!\sim\!\mathcal{U}([-\pi,\pi))^{J}$.
For the state transition we use step $\Delta t\!=\!0.1$ and process noise ${\boldsymbol{\epsilon}}_{t}\!\sim\!\mathcal{N}(\mathbf{0},Q)$ with $Q\!=\!\mathrm{diag}(0.2,\,0.2,\,10^{-2})$. We infer only the planar velocity components $\boldsymbol{\theta}=(v_{x},v_{y})$, with uniform priors $v_{x},v_{y}\sim\mathcal{U}(0.5,1.5)$. The true (data–generating) parameters are $v_{x}^{\star}\!=\!v_{y}^{\star}\!=\!1.0$ (with $v_{\phi}\!=\!0.3$ known).
Algorithmic settings.
We use an NPF with M=N=300M\!=\!N\!=\!300 particles for parameters and states (total $L\!=\!M\!\times\!N$ Monte Carlo samples per design step). We constrained the particle counts to M=NM\!=\!N and selected this value via grid search over $\{50,100,200,300,400,500\}$, choosing the smallest particle count for which performance gains became marginal relative to the added computational cost. Parameter jittering is $\kappa_{M}(\cdot\!\mid\!\boldsymbol{\theta}^{\prime})=\mathcal{N}(\boldsymbol{\theta}\!\mid\!\boldsymbol{\theta}^{\prime},\sigma_{\text{jitter}}^{2}\boldsymbol{I}_{d_{\theta}})$, with
$\sigma_{\text{jitter}}^{2}=\frac{c_{\text{jitter}}}{M^{3/2}}.$
The scale constant $c_{\text{jitter}}$ was tuned by a two-stage grid search: a coarse grid $\{0.1,0.5,1,5,10\}$, followed by a refined grid $\{0.05,0.1,0.15,0.2,0.25\}$ around the best-performing region; we set $c_{\text{jitter}}\!=\!0.15$. We resample parameter particles at every time step.
In the BAD-PODS implementation, design optimisation at each step uses stochastic gradient ascent with $K=300$ updates, where $K$ was selected via grid search over $\{50,100,200,300,400,500,600\}$. We use ADAM (Kingma and Ba, 2015) with $\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.999$, and $\varepsilon\!=\!10^{-6}$, and we tune the step size over $\alpha\in\{5\times 10^{-4},10^{-3},2.5\times 10^{-3},5\times 10^{-3},10^{-2},2\times 10^{-2}\}$, selecting $\alpha\!=\!0.01$. For the oracle implementation, we discretise the 2D design space $(-\pi,\pi]\!\times\!(-\pi,\pi]$ into a $25\!\times\!25$ grid (625 candidate values), selecting the one that maximises the EIG.
All experiments are repeated for T=50T\!=\!50 sequential design steps and averaged over 50 Monte Carlo realizations with independent random seeds.
Appendix G Ecological Growth Model
G.1 Model Description
We consider a stochastic logistic growth model with harvesting, commonly used in ecological population dynamics (Zhou et al., 2009). The latent state is the (scaled) population size $x_{t}\in\mathbb{R}_{+}$ at discrete time $t$. The design $\xi_{t}\in[0,1]$ controls harvesting intensity (e.g., fishing effort or trapping rate), and therefore affects both the transition dynamics and the observation model.
Dynamics (transition).
Let $\Delta>0$ be the integration step size. Conditioned on $(x_{t},\xi_{t})$ and parameters $\boldsymbol{\theta}=(r,k)$, the dynamics are
$x_{t+1}=x_{t}+\Delta\Bigl(r\,x_{t}\bigl(1-\tfrac{x_{t}}{k}\bigr)-q\,\xi_{t}\,x_{t}\Bigr)+\varepsilon_{t},\qquad\varepsilon_{t}\sim\mathcal{N}(0,\Delta\sigma_{x}^{2}),$ (21)
where $r>0$ is the intrinsic growth rate, $k>0$ is the carrying capacity, and $q>0$ is the catchability coefficient. The term $q\,\xi_{t}\,x_{t}$ is the (expected) harvested amount per unit time, proportional to both effort $\xi_{t}$ and available population $x_{t}$.
Observation model and design.
We define the harvested amount
$\lambda_{t}=q\,\xi_{t}\,x_{t}.$ (22)
Observations provide a noisy, partially saturated measurement of this harvested amount:
$y_{t}\!\mid\!x_{t},\xi_{t},\boldsymbol{\theta}\sim\mathcal{N}\!\bigl(h(\lambda_{t}),\,\sigma_{y}^{2}\bigr),\qquad h(\lambda_{t})=C_{\max}\,\frac{\lambda_{t}}{C_{\text{half}}+\lambda_{t}},$ (23)
where $C_{\max}>0$ is the saturation level and $C_{\text{half}}>0$ controls the transition from a linear to a saturated regime. This nonlinear response implies that increasing harvesting effort does not increase information indefinitely: for large $\lambda_{t}$ the function $h(\lambda_{t})$ saturates, obtaining virtually same outcomes for different harvesting efforts.
The design is one-dimensional, $\xi_{t}\in[0,1]$. Small $\xi_{t}$ yields weak signals $\lambda_{t}$ and therefore low signal-to-noise, while large $\xi_{t}$ can quickly depress the latent state $x_{t}$ through the transition (21) and also pushes $h(\lambda_{t})$ into a saturation regime. Together, these effects create a non-trivial trade-off and can yield intermediate optimal designs.
G.2 Simulation Setup
We simulate the growth/harvest model in (LABEL:eq:harvest-transition)–(LABEL:eq:harvest-observation) over a horizon of T=20T\!=\!20 steps, with integration step $\Delta\!=\!0.1$. The initial state is set to $x_{0}\!=\!0.4\,k$. The true (data-generating) parameters are $r^{\star}\!=\!0.5$ and $k^{\star}\!=\!300$. We assume process noise variance $\sigma_{x}^{2}\!=\!0.1$, observation noise variance $\sigma_{y}^{2}\!=\!2.0$, and saturation parameters $C_{\max}\!=\!90$ and $C_{\text{half}}\!=\!30$ are known. The design is the harvest effort $\xi_{t}\in[0,1]$, initialised at each time step as $\xi_{t}\sim\mathcal{U}(0,1)$. We infer only $\boldsymbol{\theta}\!=\!(r,k)$, with independent priors $r\sim\mathcal{U}(0.2,1.2)$ and $k\sim\mathcal{U}(200,800)$.
Algorithmic settings.
We use an NPF with M=N=200M\!=\!N\!=\!200 particles for parameters and states (total $L\!=\!M\!\times\!N$ Monte Carlo samples per design step). We constrained the particle counts to M=NM\!=\!N and selected this value via grid search over $\{50,100,200,300,400,500\}$, choosing the smallest particle count for which performance improvements became marginal relative to the added computational cost. Parameter jittering is $\kappa_{M}(\cdot\!\mid\!\boldsymbol{\theta}^{\prime})\!=\!\mathcal{N}\!\big(\boldsymbol{\theta}\!\mid\!\boldsymbol{\theta}^{\prime},\Sigma_{\text{jitter}}\big)$, with
$\Sigma_{\text{jitter}}=\begin{bmatrix}\sigma_{{r}}^{2}&0\\ 0&\sigma_{{k}}^{2}\end{bmatrix},\qquad\sigma_{{r}}^{2}=\frac{c_{{r}}}{M^{3/2}},\qquad\sigma_{{k}}^{2}=\frac{c_{{k}}}{M^{3/2}}.$
The scale constants $c_{r}$ and $c_{k}$ were tuned by a two-stage grid search: a coarse grid $\{0.1,0.5,1,5,10\}$ and $\{1,5,10,20,40,80\}$, followed by a refined grid around the best-performing region. We set $c_{r}\!=\!0.05$ and $c_{k}\!=\!50$. We resample parameter particles at every time step.
In the BAD-PODS implementation, design optimisation at each step uses stochastic gradient ascent with K=200K\!=\!200 updates, where $K$ was selected via grid search over $\{50,100,200,300,400,500,600\}$. We use ADAM (Kingma and Ba, 2015) with $\beta_{1}\!=\!0.9$, $\beta_{2}\!=\!0.999$, and $\varepsilon\!=\!10^{-6}$, and we tune the step size over $\alpha\in\{5\times 10^{-4},10^{-3},2.5\times 10^{-3},5\times 10^{-3},10^{-2},2\times 10^{-2},3\times 10^{-2},4\times 10^{-2},5\times 10^{-2}\}$, selecting $\alpha\!=\!5\times 10^{-3}$. For the oracle baseline, we discretise the 1D design space $[0,1]$ into a grid of 500 candidates and select $\xi_{t}$ by maximising the estimated EIG over this grid at each time step.
All experiments are averaged over 50 Monte Carlo realisations with independent random seeds.
Appendix H Computational Infrastructure
All experiments were conducted on a high-performance computing (HPC) cluster equipped with NVIDIA V100 GPUs. The computations made use of multi-core Intel Xeon processors, high-capacity DDR4 memory, and InfiniBand interconnects for fast data transfer between nodes. Each GPU node contained multiple V100 devices (16–32 GB memory per GPU), and all optimization and inference workloads were executed using these resources.